OpenStudy (anonymous):

find the point where there is a horizontal tangent to the equation x^3+y^3-9xy=0 other than the origin.

OpenStudy (anonymous):

These needs implicit differentiation doesn't it... D:<

OpenStudy (anonymous):

ya ;(

OpenStudy (anonymous):

not too bad

OpenStudy (anonymous):

its also asking for the vertical tangent

OpenStudy (anonymous):

ok here we go \[3x^2+3y^2y'-9y-9xy'=0\] \[3x^2-9y=9xy'-3y^2y'\] \[3x^2-9y=(9x-3y^2)y'\] \[y'=\frac{3x^2-9y}{9x-3y^2}\] \[y'=\frac{x^2-3y}{3x-y^2}\]

OpenStudy (anonymous):

check my algebra because it is late

OpenStudy (anonymous):

ya thats right

OpenStudy (anonymous):

so i need to find where that is = to 0 right?

OpenStudy (anonymous):

oh lord ok

OpenStudy (anonymous):

\[x=\pm\sqrt{3y}\]

OpenStudy (anonymous):

so i guess you replace \[x\] by \[\sqrt{3y}\] in the original equation and solve. i would use a machine

OpenStudy (anonymous):

would that give me the horiz tan?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

horizontal where the derivative is 0, this is going to be very annoying

OpenStudy (anonymous):

\[(3y)^{\frac{3}{2}}+y^3-9y^{\frac{3}{2}}=0\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

no it isn't let me fix it

OpenStudy (anonymous):

ummm idk im bad at the fraction exponets lol ill draw what i got

OpenStudy (anonymous):

\[(3y)^{\frac{3}{2}}+y^3-9\sqrt{3y}y=0\]

OpenStudy (anonymous):

\[\sqrt{3y}^3+y^3-9\sqrt{3y}y=0\] is where we start yes?

OpenStudy (anonymous):

this sucks

OpenStudy (anonymous):